# The hypothetical track-length fitting algorithm for energy measurement in liquid argon TPCs

### v2

@@ Abstract
This paper introduces the hypothetical track-length fitting algorithm, a novel method for measuring the kinetic energies of ionizing particles in liquid argon time projection chambers (LArTPCs). The algorithm finds the most probable offset in track length for a track-like object by comparing the measured ionization density as a function of position with a theoretical prediction of the energy loss as a function of the energy, including models of electron recombination and detector response. The algorithm can be used to measure the energies of particles that interact before they stop, such as charged pions that are absorbed by argon nuclei. The algorithm’s energy measurement resolutions and fractional biases are presented as functions of particle kinetic energy and number of track hits using samples of stopping secondary charged pions in data collected by the ProtoDUNE-SP detector, and also in a detailed simulation. Additional studies describe impact of the model on energy measurement performance. The method described in this paper to characterize the energy measurement performance can be repeated in any LArTPC experiment using stopping secondary charged pions.
@@ 1 Introduction
In this paper, we introduce a new method for energy measurement which can be utilized in a LArTPC. It is based on a fit for the hypothetical track-length using calorimetric and position information of hits on a reconstructed track. Data and simulation samples of the ProtoDUNE-SP detector [8, 9] are used to evaluate the performance. The method shows good performance for energy measurement on short tracks produced by interacting charged pions.
This paper is organized as follows. In Section 2, we introduce the hypothetical track-length fitting algorithm. Section 3 introduces the ProtoDUNE-SP detector and beam instrumentation, and Section 5 describes event reconstruction and the detector simulation. In Section 6, studies for the performance of the hypothetical track-length fitting algorithm are presented. Then, the paper concludes with a summary in Section 7.
2 The hypothetical track-length fitting algorithm
Before introducing the hypothetical track-length fitting algorithm, it is important to discuss the basic operational principle of a LArTPC. A particle’s activity inside liquid argon liberates electrons from argon atoms. A portion of these electrons recombines with argon ions, emitting photons. In general, a LArTPC experiment consists of a cathode plane and one or more anode planes to apply an electric field. The electrons drift toward the anode planes in this field, inducing electric signals in collection wires or pixels around the anode plane. Photon detection systems also collect the photons. The signals in the collection wires or pixels are recorded as functions of time within a time window. Since photons propagate much faster than drifting electrons, the trigger system that defines the starting point of the time window relies on photon detection systems.
@@ 2 The hypothetical track-length fitting algorithm
where , , , , and .
While the energy lost by an ionizing particle in a thin slice of material is distributed according to the Landau distribution [10], a particle traveling a long distance (more than a few ) through a material will experience an energy loss with an approximately Gaussian distribution due to the large number of ionization interactions. As a result, the continuous slowing down approximation (CSDA) based on Eq 2.1 for the average energy loss can be used to relate a particle’s initial energy to its range in the material. If a particle only undergoes multiple scattering and ionization energy loss before it stops, its initial energy can be estimated from its range with percent-level accuracy [11].
@@
The hypothetical track-length fitting algorithm is a new energy measurement method which overcomes the weaknesses of the range-based estimation and visible energy-based estimation. Figure 1 shows an example with hits of a reconstructed charged pion track. The abscissa is the residual range, which is defined for a hit on a charged particle’s trajectory as the length from the trajectory’s end point to the hit, following along the trajectory. The ordinate is in MeV/cm units. The average distribution of charged pions as a function of residual range is shown with the red solid line based on the Bethe-Bloch formula in liquid argon (LAr). A distribution of an inelastically interacting charged pion is shown with a prefit label. The measured values are obtained from the measured ionization charge density corrected for electron drift lifetime and recombination effects, as well as the charge yield expected per of energy loss in LAr. There is a gap between the prefit distribution and the vertical axis because first and last several hits are not included in the fit. They could be hits with incomplete or only noise. This is discussed in more detail in Section 6.2. To measure the energy of a charged pion that has interacted inelastically with the argon, we look for an offset in the residual range which makes the residual range and distribution agree best with the known physics for passage of charged particle through liquid argon. That offset is added to the original track length to estimate the hypothetical track-length. This is the track length that the particle with the prefit distribution would have traveled until it stopped if it did not interact inelastically. Finally, energy at the starting point of the track is measured using the CSDA and the hypothetical track-length. The post fit distribution with the best offset is also shown.
Based on the residual range vs. curve derived from the Bethe-Bloch formula, shown in figure 1, it could be easily expected that energy measurement performance will highly depend on particle’s initial kinetic energy (). Short residual range region that corresponds to low of charged pions exhibits stiff slopes, commonly referred to as the Bragg peak region. Since varies rapidly as a function of residual range, it is easier to find the best offset in this region compared to longer residual range region, known as the MIP (minimum-ionizing particle) region, that has much smaller slopes. For an inelastically interacting charged pion track for which all hits have MIP-like energy deposit around 2 , the hypothetical track-length fitting algorithm cannot work well. This is shown in Section 6.
One of the most important parts of the hypothetical track-length fitting is the procedure to calculate the best offset. This paper describes and compares two methods. One method is based on minimizing a function assuming Gaussian distributions of inputs, and the second method is based on maximizing a likelihood function.
@@ 2.1 Gaussian approximation
The simplest way to score the agreement between the Bethe–Bloch curve and the measured distribution is by calculating the between them. For a given position offset ,
@@ 2.2 Maximum-likelihood
Using the PDFs where stands for the probability, a joint likelihood () is calculated by multiplying the probabilities for each hit. For hypothetical track-length fitting, the expected of particle at a hit is calculated using the CSDA and the length given by sum of hit’s residual range and an additional offset. As a result, the total likelihood for a given offset, , is written as
@@ 5 Simulation and reconstruction
In this paper, data events with incident beam particles inside ProtoDUNE-SP are used to study the performance of the hypothetical track-length fitting method.
@@ 6 Energy measurement performance
The performance of the hypothetical track-length fitting method is tested using data and MC simulation samples of the ProtoDUNE-SP experiment. For this purpose, stopping charged pions are selected from the secondary particles resulting from the interaction between the charged pion beam and argon atoms. So, we introduce the selection for stopping charged pions first. It is validated that the CSDA can describe well the true for selected stopping charged pion candidates. As a result, from the CSDA is used as reference to compare performance of the hypothetical track-length fitting method between data and MC simulation samples. Tunable parameters of the hypothetical track-length fitting algorithm are also introduced. Then, measurement performance of the hypothetical track-length fitting algorithm is presented. The fitting is performed using subsets of stopping charged pions’ hits to mimic inelastically interacting charged pions. Therefore, algorithm performance is presented as functions of and number of hits that are used for the fitting. For the MC simulation sample, the energy measurement performance is also presented using true as reference with the same selection that is used for comparison between data and MC simulation samples. Finally, a systematic study related with impact of modeling is performed to understand energy measurement scale difference between data and MC simulation samples.
@@ 6.1 Stopping charged pions as a validation sample
Here, is the expected for a given residual range coming from the CSDA using the Bethe-Bloch formula in Eq. 2.1, and contains uncertainties of for both from the CSDA and the TPC’s energy measurement resolution. Since Eq. 6.1 uses hits with residual range less than 26 cm to make the not to be dominated by MIP hits, stopping charge pions with visible Bragg peaks have smaller values compared to inelastically interacting charged pions. Figure 5 shows the distribution as a function of true in a MC sample of charged pions. A peak with value about from 2 to 4 is coming from stopping charged pions. The other peak that has value around 12 to 16 is coming from inelastically interacting charged pions that left tracks with length almost 26 cm or longer where all hits are MIP-like. That is the reason why the distribution of the band is approximately constant as a function of true in the two dimensional distribution. Very high values greater than 20 are coming from inelastically interacted charged pions that left short tracks. In this study, value for the secondary charged pions is required to be smaller than 6.0 to select the stopping charged pions. Figure 6 shows that the CSDA with the full track length () estimates the true () well after applying the cut.
@@ 6.2 Algorithm parameters
The first parameter is the minimum number of usable hits to include in the fit. In the case where two PDFs overlap significantly between two different values of , including only a small number of hits in the fit can lead to a huge uncertainty in the energy measurement. For example, the PDFs of charged pions with 200 MeV and 500 MeV of values have similar shapes and a large overlap as shown in figure 3. Therefore, fitting a small number of hits which have random values given by PDFs will perform poorly. On the other hand, requiring many hits excludes low-energy particles, since is highly correlated with the total length of the particle’s trajectory. In this paper, at least 15 hits are required. This number corresponds to approximately 40 MeV of kinetic energy threshold for charged pions inside the ProtoDUNE-SP TPC. This requirement can be optimized depending on the targeted range of the particles under study.
@@ 6.3 Performance in MC simulation sample using true as the reference
The CSDA using an incomplete track () underestimates the charged pion’s energy as shown in figure 7. The hypothetical track-length fitting algorithm can provide better energy measurement () by introducing best additional track length. Figure 8 shows two-dimensional distributions of the fractional energy residual from the Gaussian approximation and the maximum-likelihood methods. The Gaussian approximation shows clear bands corresponding to biased energy measurements. These measurements occur when the best additional track length is found to be in the minimum ionizing .
@@
This bias implies that the maximum-likelihood method is the most appropriate one for charged pions. Therefore, we fix the fitting method to maximum-likelihood. Figure 9 shows examples of histograms which are used to measure resolutions and fractional biases. The energy measurement resolution and fractional bias for each distribution are calculated using data points within of the fitted Gaussian function. Figure 10 shows a summary of extracted resolutions and fractional biases. The resolutions are better than 6% and the fractional biases are smaller than 5%.
@@ 6.4 Performance in data and MC simulation samples using as the reference
Figure 14 summarizes the energy measurement performance for a pure sample of secondary charged pions from the MC simulation. Figure 15 shows a direct performance comparison between data and MC simulation samples. The upper cutoff for each number of hits range is lower than that in figure 14 to ensure good modeling of the proton contribution. The top plot of figure 15 shows good agreement between MC and data in resolution with values smaller than 8%. The bottom plot of figure 15 shows a fractional bias result that is smaller than 10% with discrepancy between MC simulation and data. It means that there should be studies of the energy scale correction and the corresponding systematic uncertainty in order to utilize the hypothetical track-length fitting method. The energy scale correction can be measured at any LArTPC using stopping secondary charged pions from neutrino interactions as introduced in this section, but as a function of rather than the . Section 6.5 describes the corresponding systematic uncertainty estimates, separately for each source.
An additional comment is that the hypothetical track-length fitting method does not work well for a track for which the values have yet to sample the Bragg peak. It is a major reason why there is an upper cutoff for each number of hits range. It is an energy measurement method mainly for charged pions with less than 400 MeV. In this energy region, the most dominant inelastic scattering topology between charged pions and argon nuclei is absorption.
@@ 6.5 Impact of modeling
Small discrepancies are seen between data and MC simulation, and these motivate ionization scale and resolution corrections with corresponding systematic uncertainties. Estimations of these corrections are needed in order to use the hypothetical track-length fitting method for physics analyses. In this section, we present the impact of modeling on results in figure 15 to understand the discrepancy between MC and data in the fractional bias.
@@
Recombination between argon ions and electrons has a significant impact on conversion between collected electric charge per unit length [electrons/cm] and [MeV/cm]. Thus, it should be modeled.
The ProtoDUNE-SP experiment uses the modified box model [27] to consider electron recombination. In this model, the relation between and is
@@
As a result, figure 16-(d) shows good agreement between data and MC simulation for distributions of beam muon hits in a residual range region from 95 to 96 cm, which corresponds to the MIP region. But, for hits near the Bragg peak region, the agreement between data and MC simulation deteriorates as shown in the plots with shorter residual ranges of figure 16. It means that the modified box model with parameters in [27] cannot describe the recombination effect in ProtoDUNE-SP in the MIP region and the Bragg peak region at the same time. It is well illustrated in figure 17. The abscissa shows fitted MPVs of MC simulation sample for beam muon hits in residual range region from 2 to 100 with 1 step. The ordinate shows the ratio of fitted MPV between MC simulation and data in each residual range region. MPV ratio is close to unity where MC simulation MPV is smaller than 1.7 . With increasing MC simulation MPV values, the ratio goes down to about 0.985 and goes up to about 1.04. To understand the impact of such differences in the probability density function between data and MC simulation on fractional bias results shown in figure 15, we performed studies described below.

### v3

# The track-length extension fitting algorithm for energy measurement of interacting particles in liquid argon TPCs and its performance with ProtoDUNE-SP data

@@ Abstract
This paper introduces a novel track-length extension fitting algorithm for measuring the kinetic energies of inelastically interacting particles in liquid argon time projection chambers (LArTPCs). The algorithm finds the most probable offset in track length for a track-like object by comparing the measured ionization density as a function of position with a theoretical prediction of the energy loss as a function of the energy, including models of electron recombination and detector response. The algorithm can be used to measure the energies of particles that interact before they stop, such as charged pions that are absorbed by argon nuclei. The algorithm’s energy measurement resolutions and fractional biases are presented as functions of particle kinetic energy and number of track hits using samples of stopping secondary charged pions in data collected by the ProtoDUNE-SP detector, and also in a detailed simulation. Additional studies describe the impact of the model on energy measurement performance. The method described in this paper to characterize the energy measurement performance can be repeated in any LArTPC experiment using stopping secondary charged pions.
@@ 1 Introduction
In this paper, we introduce a new method for energy measurement which can be utilized in a LArTPC. It is based on a fit for the best track-length extension using calorimetric and position information of hits on a reconstructed track. For an inelastically interacting particle, the method calculates the total track length the particle would have traveled if it had not interacted inelastically. Therefore, the method addresses the challenges introduced in the above paragraph. Data and simulation samples of the ProtoDUNE-SP detector [8, 9] are used to evaluate the performance. The method shows good performance for energy measurement on short tracks produced by interacting charged pions.
This paper is organized as follows. In Section 2, we introduce the track-length extension fitting (TLEFit) algorithm. Section 3 introduces the ProtoDUNE-SP detector and beam instrumentation, and Section 5 describes event reconstruction and the detector simulation. In Section 6, studies for the performance of the TLEFit algorithm are presented. Then, the paper concludes with a summary in Section 7.
2 The track-length extension fitting algorithm
Before introducing the track-length extension fitting (TLEFit) algorithm, it is important to discuss the basic operational principle of a LArTPC. A particle’s activity inside liquid argon liberates electrons from argon atoms. A portion of these electrons recombines with argon ions, emitting photons. In general, a LArTPC experiment consists of a cathode plane and one or more anode planes to apply an electric field. The electrons drift toward the anode planes in this field, inducing electric signals in collection wires or pixels around the anode plane. Photon detection systems also collect the photons. The signals in the collection wires or pixels are recorded as functions of time within a time window. Since photons propagate much faster than drifting electrons, the trigger system that defines the starting point of the time window relies on photon detection systems.
@@ 2 The track-length extension fitting algorithm
where , , , , and . For kinetic energies above tens of MeV, which is the threshold for particles to be reconstructed as tracks, nuclear effects in energy deposition is negligible.
While the energy lost by an ionizing particle in a thin slice of material is distributed according to the Landau distribution [10], a particle traveling a long distance (more than a few ) through a material will experience an energy loss with an approximately Gaussian distribution due to the large number of ionization interactions and the central limit theorem. As a result, the continuous slowing down approximation (CSDA) based on Eq 2.1 for the average energy loss can be used to relate a particle’s initial energy to its range in the material [10]. If a particle only undergoes multiple scattering and ionization energy loss before it stops, its initial energy can be estimated from its range with percent-level accuracy [11].
@@
The TLEFit algorithm introduced in this paper is a new energy measurement method which overcomes the weaknesses of the range-based estimation and visible energy-based estimation. Figure 1 shows an example with hits of a reconstructed charged pion track. The abscissa is the residual range, which is defined for a hit on a charged particle’s trajectory as the length from the trajectory’s end point to the hit, following along the trajectory. The ordinate is in MeV/cm units. The average distribution of charged pions as a function of residual range is shown with the red solid line based on the Bethe-Bloch formula in liquid argon (LAr). A distribution of an inelastically interacting charged pion is shown with a prefit label. The measured values are obtained from the measured ionization charge density corrected for electron drift lifetime and recombination effects, as well as the charge yield expected per of energy loss in LAr. There is a gap between the prefit distribution and the vertical axis because first and last several hits are not included in the fit. They could be hits with incomplete or only noise. This is discussed in more detail in Section 6.2. To measure the energy of a charged pion that has interacted inelastically with the argon, we look for an offset in the residual range which makes the residual range and distribution agree best with the known physics for passage of charged particle through liquid argon. That offset is added to the original track length to estimate the best track-length. This is the track length that the particle with the prefit distribution would have traveled until it stopped if it did not interact inelastically. Finally, energy at the starting point of the track is measured using the CSDA and the best track-length. The post fit distribution with the best offset is also shown.
Based on the residual range vs. curve derived from the Bethe-Bloch formula, shown in figure 1, it could be easily expected that energy measurement performance will highly depend on particle’s initial kinetic energy (). Short residual range region that corresponds to low of charged pions exhibits stiff slopes, commonly referred to as the Bragg peak region. Since varies rapidly as a function of residual range, it is easier to find the best offset in this region compared to longer residual range region, known as the MIP (minimum-ionizing particle) region, that has much smaller slopes. For an inelastically interacting charged pion track for which all hits have MIP-like energy deposit around 2 , the TLEFit algorithm cannot work well. This is shown in Section 6.
One of the most important parts of the TLEFit is the procedure to calculate the best offset. This paper describes and compares two methods. One method is based on minimizing a function assuming Gaussian distributions of inputs, and the second method is based on maximizing a likelihood function.
@@ 2.1 Gaussian approximation
The simplest way to score the agreement between the Bethe–Bloch curve and the measured distribution is by calculating the between them. Many LArTPC experiments have employed -based particle identification scores [14]. Therefore, we describe this method before introducing a likelihood-based method.
For a given position offset , we define
@@ 2.2 Maximum-likelihood
Using the PDFs where stands for the probability, a joint likelihood () is calculated by multiplying the probabilities for each hit. For the TLEFit, the expected of particle at a hit is calculated using the CSDA and the length given by sum of hit’s residual range and an additional offset. As a result, the total likelihood for a given offset, , is written as
@@ 5 Simulation and reconstruction
In this paper, data events with incident beam particles inside ProtoDUNE-SP are used to study the performance of the TLEFit method.
@@ 6 Energy measurement performance
The performance of the TLEFit method is tested using data and MC simulation samples of the ProtoDUNE-SP experiment. For this purpose, stopping charged pions are selected from the secondary particles resulting from the interaction between the charged pion beam and argon atoms. So, we introduce the selection for stopping charged pions first. It is validated that the CSDA can describe well the true for selected stopping charged pion candidates. As a result, from the CSDA is used as reference to compare performance of the TLEFit method between data and MC simulation samples. Tunable parameters of the TLEFit algorithm are also introduced. Then, measurement performance of the TLEFit algorithm is presented. The fitting is performed using subsets of stopping charged pions’ hits to mimic inelastically interacting charged pions. Therefore, algorithm performance is presented as functions of and number of hits that are used for the fitting. For the MC simulation sample, the energy measurement performance is also presented using true as reference with the same selection that is used for comparison between data and MC simulation samples. Finally, a systematic study related with impact of modeling is performed to understand energy measurement scale difference between data and MC simulation samples.
@@ 6.1 Stopping charged pions as a validation sample
Here, is the expected for a given residual range coming from the CSDA using the Bethe-Bloch formula in Eq. 2.1, and contains uncertainties of for both from the CSDA and the TPC’s energy measurement resolution. Since Eq. 6.1 uses hits with residual range less than 26 cm to make the not to be dominated by MIP hits, stopping charge pions with visible Bragg peaks have smaller values compared to inelastically interacting charged pions. To produce a normalized , number of hits included into the , , is divided. Figure 5 shows the distribution as a function of true in a MC sample of charged pions. A peak with value about from 2 to 4 is coming from stopping charged pions. The other peak that has value around 12 to 16 is coming from inelastically interacting charged pions that left tracks with length almost 26 cm or longer where all hits are MIP-like. That is the reason why the distribution of the band is approximately constant as a function of true in the two dimensional distribution. Very high values greater than 20 are coming from inelastically interacted charged pions that left short tracks. In this study, value for the secondary charged pions is required to be smaller than 6.0 to select the stopping charged pions. Figure 6 shows that the CSDA with the full track length () estimates the true () well after applying the cut.
@@ 6.2 Algorithm parameters
The first parameter is the minimum number of usable hits to include in the fit. In the case where two PDFs overlap significantly between two different values of , including only a small number of hits in the fit can lead to a huge uncertainty in the energy measurement. For example, the PDFs of charged pions with 200 MeV and 500 MeV of values have a large overlap as shown in figure 3. Therefore, fitting a small number of hits which have random values given by PDFs will perform poorly. On the other hand, requiring many hits excludes low-energy particles, since is highly correlated with the total length of the particle’s trajectory. In this paper, at least 15 hits are required. This number corresponds to approximately 40 MeV of kinetic energy threshold for charged pions inside the ProtoDUNE-SP TPC. This requirement can be optimized depending on the targeted range of the particles under study.
@@ 6.3 Performance in MC simulation sample using true as the reference
The CSDA using an incomplete track () underestimates the charged pion’s energy as shown in figure 7. The TLEFit algorithm can provide better energy measurement () by introducing the best additional track length. Figure 8 shows two-dimensional distributions of the fractional energy residual from the Gaussian approximation and the maximum-likelihood methods. The Gaussian approximation shows clear bands corresponding to biased energy measurements. These measurements occur when the best additional track length is found to be in the minimum ionizing . The maximum-likelihood method plots show populations with biased energy measurements, but these are not concentrated into a band, as seen in the Gaussian approximation plots.
@@
This bias implies that the maximum-likelihood method is the most appropriate one for charged pions. Therefore, we fix the fitting method to maximum-likelihood. Figure 9 shows examples of histograms which are used to measure resolutions and fractional biases. The energy measurement resolution and fractional bias for each distribution are calculated using data points within of the fitted Gaussian function.
Figure 10 summarizes the extracted resolutions and fractional biases. The resolutions are better than 6.5%, and the fractional biases are smaller than 4%. Better resolutions are observed at lower and with a greater number of hits. Lower benefits from the steeper slope of as a function of , while a greater number of hits provides stronger constraints. Sizes of fractional biases are smaller than 1.5%, except for charged pions with 15 to 30 hits, which exhibit biases down to -3.5% at lower .
From a study using MC truth information, we find that tracks with fewer hits tend to have shorter reconstructed total track lengths compared to the true track lengths. For a given , a smaller number of hits implies that the track’s angle relative to the anode plane is steeper, reducing the number of sense wires that can collect signals from the track. In such cases, the 3D spatial points of hits are reconstructed with larger uncertainties. Consequently, the residual ranges and their sum, the reconstructed total track length, are biased to be shorter than the true values. This effect is most pronounced in the smallest-hit region. However, the magnitude of the fractional bias, up to 3.5%, remains relatively small compared to the biases observed when using CSDA for inelastically interacting charged pions, as shown in Figure 7.
@@ 6.4 Performance in data and MC simulation samples using as the reference
Figure 14 summarizes the energy measurement performance for a pure sample of secondary charged pions from the MC simulation. Figure 15 shows a direct performance comparison between data and MC simulation samples. The upper cutoff for each number of hits range is lower than that in figure 14 to ensure good modeling of the proton contribution. The top plot of figure 15 shows good agreement between MC and data in resolution with values smaller than 8%. The bottom plot of figure 15 shows a fractional bias result that is smaller than 10% with discrepancy between MC simulation and data. It means that there should be studies of the energy scale correction and the corresponding systematic uncertainty in order to utilize the TLEFit method. The energy scale correction can be measured at any LArTPC using stopping secondary charged pions from neutrino interactions as introduced in this section, but as a function of rather than the . Section 6.5 describes the corresponding systematic uncertainty estimates, separately for each source.
An additional comment is that the TLEFit method does not work well for a track for which the values have yet to sample the Bragg peak. It is a major reason why there is an upper cutoff for each number of hits range. It is an energy measurement method mainly for charged pions with less than 400 MeV. In this energy region, the most dominant inelastic scattering topology between charged pions and argon nuclei is absorption.
@@ 6.5 Impact of modeling
Small discrepancies are seen between data and MC simulation, and these motivate ionization scale and resolution corrections with corresponding systematic uncertainties. Estimations of these corrections are needed in order to use the TLEFit method for physics analyses. In this section, we present the impact of modeling on results in figure 15 to understand the discrepancy between MC and data in the fractional bias. Motivation of testing modeling is that data shows up to about 7% fractional bias with respect to the range-based energy using full track length while MC sample has less than 2.5% fractional bias. We observe that modeling could explain the discrepancy.
@@
Recombination between argon ions and electrons has a significant impact on conversion between collected electric charge per unit length [electrons/cm] and [MeV/cm]. Thus, it should be modeled. The ProtoDUNE-SP experiment uses the modified box model [28] to consider electron recombination. In this model, the relation between and is
@@
As a result, figure 16-(d) shows good agreement between data and MC simulation for distributions of beam muon hits in a residual range region from 95 to 96 cm, which corresponds to the MIP region. But, for hits near the Bragg peak region, the agreement between data and MC simulation deteriorates as shown in the plots with shorter residual ranges of figure 16. It means that the modified box model with parameters in [28] cannot describe the recombination effect in ProtoDUNE-SP in the MIP region and the Bragg peak region at the same time. It is well illustrated in figure 17(a). The abscissa shows fitted MPVs of MC simulation sample for beam muon hits in residual range region from 2 to 100 with 1 step. The ordinate shows the ratio of fitted MPV between MC simulation and data in each residual range region. MPV ratio is close to unity where MC simulation MPV is smaller than 1.7 . With increasing MC simulation MPV values, the ratio goes down to about 0.985 and goes up to about 1.04. To understand the impact of such differences in the probability density function between data and MC simulation on fractional bias results shown in figure 15, we performed studies described below.